# Special Hamiltonian $S^{1}$-actions on symplectic 4-manifolds

### v3

@@ 3 Some topological constraints imposed by Maslov condition
(ii) It follows from (i) that an equivariant version of Darboux Theorem [1] (see also [6] Theorem 22.2) applies to every isolated fixed point $p\in Fix(G)$. Namely, for each isolated fixed point $p\in Fix(G)$, there exists a $G$-invariant neighborhood $U$ of $p\in M$ and a $G$-equivariant diffeomorphism $f$ of $U$ into $(M,\omega)$ with $f(p)=p$, $f^{*}\omega=\omega_{0}$ is the standard symplectic form $\omega_{0}=\sum_{k=1}^{2}dx_{k}\wedge dy_{k}$ on $\mathbb{R}^{4}$, and $f^{-1}\circ G\circ f=\textsf{G}\subset SU(2)$. A proof of (ii) is given below.
Let $p\in Fix(G)$ be an isolated fixed point. By Darboux theorem we may identify $\omega$ with the standard symplectic form on $\omega_{0}:=\sum_{j=1}^{2}dx_{j}\wedge dy_{j}$ on $\mathbb{R}^{4}$ when restricted to some Darboux chart $U\subset M$ centered at $p$, on which $\omega=\omega_{0}$ is $G$-invariant.
For simplicity we assume that the $G$-action near $p$ is semi-free with $m_{p}=1$. Then the linear special Hamiltonian $S^{1}$ action $G^{L}=G_{*}:T_{p}M\to T_{p}M$ is
$G^{L}=\Big{\{}G^{L}_{\theta}:=\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix}\Big{|}\ \theta\in\mathbb{R}/2\pi\mathbb{Z}\Big{\}}=\textsf{G}\subset SU(2)$
up to conjugation. Identify $(T_{p}M,\omega_{p})$ with $(\mathbb{R}^{4},\omega_{0})$ and let $g_{0}$ denote the standard Euclidean metric on $(T_{p}M,\omega_{p})\simeq(\mathbb{R}^{4},\omega_{0})$. $g_{0}$ is $\omega_{0}$-compatible and is invariant under the G-action on $\mathbb{R}^{4}$. Also let $g$ denote a $\omega$-compatible Riemannian metric on $M$ which is also $G$-invariant at least on a $G$-invariant compact domain $D\subset M$ with $p$ as its central interior point, and $D\cap Fix(G)=\{p\}$ We may assume that $g=g_{0}$ when restricted to $T_{p}M$.
Consider the exponential map with respect to the Riemannian metric $g$ on $M$:
$\EuFrak{e}_{p}:=\exp_{p}:T_{p}M\cong\mathbb{R}^{4}\to M,\quad\EuFrak{e}_{p}(v):=\gamma_{v}(1)$
where $\gamma_{v}(t)$, $t\in\mathbb{R}$, is the geodesic emanated from $p$ with $\gamma_{v}(0)=p$, $\dot{\gamma}(0)=v$. There exists $\epsilon>0$ such that
$\EuFrak{e}_{p}:\ V:=\{v\in T_{p}M\mid|v|<\epsilon\}\to\EuFrak{e}_{p}(v)=\gamma_{v}(1)$
is a diffeomorphism from $V$ onto $U:=\EuFrak{e}_{p}(V)\subset M$. For $v_{0}\in V\subset T_{p}M$, $v_{0}\neq 0$, let $v_{\theta}:=G^{L}_{\theta}(v_{0})$, $|v_{\theta}|=|v_{0}|$ with respect to both $g_{0}$ and $g$. Since $g$ is $G$-invariant on $U$, the $G$-action preserves the space of $g$-geodesics $\gamma_{v}$, $v\in V$, with $\gamma_{v_{\theta}}(t)=\gamma_{tv_{\theta}}(1)=G_{\theta}(\gamma_{tv_{0}}(1))=G_{\theta}(\gamma_{v_{0}}(t))=G_{\theta}(\gamma_{tv_{0}}(1))$, $|t|\leq 1$, $\theta\in\mathbb{R}/2\pi\mathbb{Z}$.
Consider $\EuFrak{e}^{-1}_{p}\circ G\circ\EuFrak{e}_{p}:V\to V$. At $v_{0}\in V$,
$\EuFrak{e}^{-1}_{p}\circ G_{\theta}\circ\EuFrak{e}_{p}(v_{0})=\EuFrak{e}^{-1}_{p}\circ G_{\theta}\circ\gamma_{v_{0}}(1)=\EuFrak{e}^{-1}_{p}(\gamma_{v_{\theta}}(1))=v_{\theta}=G^{L}_{\theta}(v),$
$\Rightarrow\quad\EuFrak{e}^{-1}_{p}\circ G\circ\EuFrak{e}_{p}=\textsf{G}\quad\text{ on }V$
or, equivalently, $G=\EuFrak{e}_{p}\circ\textsf{G}\circ\EuFrak{e}_{p}^{-1}$ on $U$. Also let $\EuFrak{e}_{p}^{0}$ denote the exponential map with respect to $g_{0}$.
WLOG we may identify $(V,\omega_{0})$ with $(U_{0}:=\EuFrak{e}_{p}^{0}(V),\omega_{0})$. Then $\EuFrak{e}_{p}:U_{0}\to U$ is a diffeomorphism from $(U_{0},\omega_{0})$ onto $(U,\omega_{0})$. Let
$\omega_{1}:=(\EuFrak{e}_{p}^{-1})^{*}\omega_{0}=\sum_{j=1}^{2}x^{\prime}_{j}\wedge dy^{\prime}_{j},\quad x^{\prime}_{j}=x_{j}\circ\EuFrak{e}_{p}^{-1},\quad y^{\prime}_{j}=y_{j}\circ\EuFrak{e}_{p}^{-1}.$
Up to the difffeomorphism $\EuFrak{e}_{p}:(U_{0},\omega_{0})\to(U,\omega_{0})$ $G$ acts on $U$ symplectically and linearly with respect to the symplectic form $\omega_{1}$, with symplectic basis $\{\frac{\partial}{\partial x^{\prime}_{1}},\ \frac{\partial}{\partial y^{\prime}_{1}},\ \frac{\partial}{\partial x^{\prime}_{2}},\ \frac{\partial}{\partial y^{\prime}_{2}}\}$. Now
$G=\EuFrak{e}_{p}\circ\textsf{G}\circ\EuFrak{e}^{-1}_{p},$
[…]
$(X:=(\EuFrak{e}_{p})_{*}\textsf{X},\ h:=\textsf{h}\circ\EuFrak{e}^{-1}_{p})$
are the the corresponding Hamiltonian vector field and Hamiltonian function associated to the $G$-action on $U$.
We may extend $\EuFrak{e}_{p}$ to a diffeomorphism $\EuFrak{e}^{\prime}_{p}:M\to M$ from $M$ onto $M$ with $G$-invariant compact support containing $U_{0}\cup U$. Let $G^{\prime}:=\EuFrak{e}^{\prime}_{p}\circ\textsf{G}\circ(\EuFrak{e}_{p}^{\prime})^{-1}$. $G^{\prime}$ acts as a Hamiltonian $S^{1}$-action $M$, with $G^{\prime}=G$ on $U$ and outside the compact support of $\EuFrak{e}^{\prime}_{p}$.
In addition, we claim that there exists a smooth isotopy $\phi_{t}:U\to M$, $t\in[0,1]$, with $\phi_{0}=id$, $\phi_{t}(p)=p$ for all $t$, and $\phi_{1}^{*}\omega_{1}=\omega_{0}$.
Let $\omega_{t}:=(1-t)\omega_{0}+t\omega_{1}$, $t\in[0,1]$, then $\omega_{t}|_{p}=\omega_{0}|_{p}$ for all $t$. By shrinking $U$ if necessary we may assume that $\omega_{t}$ are symplectic on $U$ for all $t\in[0,1]$. Since the $G$-action is special Hamiltonian, $\omega_{t}$ are exact for $t\in[0,1]$,
$\omega_{t}=d\alpha_{t},\quad\frac{d\omega_{t}}{dt}=\omega_{1}-\omega_{0}=d(\alpha_{1}-\alpha_{0})$
for some $\alpha_{t}\in\Omega^{1}(M)$. The $G$-action preserves both $\omega_{1}$ and $\omega_{0}$ on $U$, hence all $\omega_{t}$ are $G$-invariant. Since $G$ is compact, by averaging we may assume that all $\alpha_{t}$ are $G$-invariant at least on $U$ by shrinking $U$ if necessary.
Assume there exists an isotopy $\phi_{t}$, $t\in[0,1]$, $\phi_{0}=id$, such that $\phi^{*}_{t}\omega_{t}=\omega_{0}$, and let $X_{t}$ be the vector field associated to $\phi_{t}$, then
$0=\frac{d}{dt}\Big{(}\phi^{*}_{t}\omega_{t}\Big{)}=\phi^{*}_{t}\Big{(}\frac{d\omega_{t}}{dt}+L_{X_{t}}\omega_{t}\Big{)}=\phi^{*}_{t}d\Big{(}\alpha_{1}-\alpha_{0}+\omega_{t}(X_{t}.\cdot))\Big{)}$
$\Rightarrow\omega_{t}(X_{t},\cdot)=\alpha-\alpha_{0}$ (2)
As $\omega_{t}$’s are symplectic on $U$, for each $t\in[0,1]$ there exists a unique vector field $X_{t}$ defined on $U$ with (2) satisfied . Both $\omega_{t}$ and $\alpha_{t}$ are $G$-invariant on $U$, so are $X_{t}$. Moreover, $X_{t}=0$ at $p$ since $\omega_{t}=\omega_{0}$ at $p$ for $t\in[0,1]$. Then $\phi_{t}:U\to M$ are defined and equivariant with respect to $G$ on some $G$-invariant neighborhood $U_{p}\subset U$ of $p$, with $\phi_{0}=id$, $\phi_{1}^{*}\omega_{1}=\omega_{0}$, $\phi_{1}\circ\textsf{G}\circ\phi_{1}^{-1}=G$. This finishes the proof of (ii).
@@
For (iii), observe that $dh|_{p}=-\omega(X_{h}|_{p},\cdot)$ for $p\in M$. Since $\omega$ is nondegenerate, $\omega(X_{h}|_{p},\cdot)=0$ iff $X_{h}$ vanishes at $p$, i.e., iff $p\in Fix(G)$, i.e., $Crit(h)=Fix(G)$. By (ii) we may assume that $=\textsf{G}$ G and hence $=\textsf{h}$ h near each isolated critical point $p\in Fix(G)$ of $h$, therefore all isolated critical points of $h$ are nondegenerate with Morse index $2$. ∎
@@
Forgetting the Maslov condition, 4-dimensional special Hamiltonian $S^{1}$-spaces are by themselves (not necessarily compact) Hamiltonian $S^{1}$-spaces. Recall that 4-dimensional compact Hamiltonian $S^{1}$-spaces have been classified by Karshon [8]:
@@
[…]
@@
Recall that if $(M,\omega,G)$ is a 4-dimensional compact Hamiltonian $S^{1}$-space then each connected component of $Fix(G)$ is either a single point or a symplectic surface, and the maximum and minimum of the moment map $h$ is each attained on exactly one component of $Fix(G)$ (see e.g. [8] Appendix A).
@@
A closed compact symplectic 4-manifold $(M,\omega)$ does not admit a special Hamiltonian $S^{1}$-action.
@@
Assume in contrary that $(M,\omega)$ allows a special Hamiltonian $S^{1}$-action $G:S^{1}\to Ham(M,\omega)$. We may assume that the action is effective. Let $h$ denote its moment map. Then $Fix(G)=Crit(h)$. $M$ is compact, $h$ is bounded, and the sets $M_{-}:=\{p\in M\mid h(p)=\min h\}$ and $M_{+}:=\{p\in M\mid h(p)=\max h\}$ are both nonempty. Since every isolated fixed point of the $G$-action has Morse index 2, the set $M_{-}$ contains no connected component with isolated fixed points, and neither does $M_{+}$. Then both $M_{-}$ and $M_{+}$ are connected compact closed symplectic surfaces and, following [8], $M$ can be obtained from a ruled manifold with two fixed surfaces and no interior fixed points by a sequence of equivalent blowups at fixed points that are not minima for the moment map of the ruled manifold.
Suppose that $M$ is a ruled manifold, then the orbit $\gamma_{q}:=Orb_{G}(q)$ of $q\not\in Fix(G)$ is contained some symplectic $S^{2}$-fiber $S$ of $M$. But $c_{1}(M)([S])=S\cdot S+e(S)=0+2\neq 0$, $c_{1}(M.\omega)\neq 0$. Indeed, let $F_{p}$ denote the $S^{2}$ fiber over $p\in M_{-}$. Fix any $q\in F_{p}\setminus(M_{+}\cup M_{-})$. $Orb_{G}(q)$ divides $F_{p}\cong S^{2}$ into two discs $D_{-}\cup D_{+}$ with $D_{-}\cap M_{-}=\{p\}$. We may identity a $G$-invariant symplectic Darboux neighborhood of $p\in M$ as $D^{\prime}\times D_{\epsilon}\subset\mathbb{C}_{z_{1}}\times\mathbb{C}_{z_{2}}$ so that $D_{-}\subset D^{\prime}\times\{0\}$, and $D_{\epsilon}$ is a coordinate chart of $F_{p}$ centered at $p$. Up to a choice of orientation of $\gamma_{q}$ we may assume that the tangent vector field along $\gamma_{q}$ is $X:=-y_{1}\partial_{x_{1}}+x_{1}\partial_{y_{1}}$. Let $v:=\partial_{x_{2}}|_{\gamma_{q}}$ be the vector field $\partial_{x_{2}}$ restricted to $\gamma_{q}$. Then $L:=X\wedge v$ is a Lagrangian plane field along $\gamma_{q}$ with Maslov index $\mu(L,D_{-})=2$. On the other hand, if we choose $D_{+}$ as the disc bounded by $\gamma_{q}$, then the orientation of $\gamma_{q}$ as the boundary of $D_{+}$ is given by $-X$, and the corresponding Maslov index becomes $\mu(L=(-X)\wedge v,D_{+})=-2$. Either way the Maslov condition is not satisfied. So $M$ cannot be a ruled manifold, nor can it be obtained from any ruled manifold by a sequence of equivalent blowups at fixed points that are not minima for the moment map of the ruled manifold. We conclude that a 4-dimensional compact symplectic manifold $(M,\omega)$ does not admit an effective special Hamiltonian $S^{1}$-action. ∎
@@
Let $(M,\omega)$ be a connected noncompact symplectic 4-manifold. Suppose that there is a special Hamiltonian $S^{1}$-action $G:S^{1}\to Ham(M,\omega)$ on $M$ with contractible orbits and nonempty fixed point set $Fix(G)$, then $Fix(G)$ consists of isolated points.
@@
Let $G:S^{1}\to Ham(M,\omega)$ be a special Hamiltonian $S^{1}$-action on $(M,\omega)$. Denote by $X$ the Hamiltonian vector field on $M$ which generates the $G$-action, and $h:M\to\mathbb{R}$ the corresponding moment map defined by $\omega(X,\cdot)=-dh$. By assumption all $G$-orbits are contractable and $Fix(G)$ is not empty. It is known that every connected component of $Fix(G)$ is either a single point or a symplectic surface ([8], Appendix A). If $Fix(G)$ contains a connected symplectic surface $\Sigma$ then $h$ is constant on $\Sigma$ and $h(\Sigma)$ is a local extremum of $h$.
Fix a point $p\in\Sigma$ and an open neighborhood $U_{p}\subset\Sigma$ of $p$ so that $G$ acts nontrivially near $U_{p}$ except on $U_{p}$, and with isomorphic stabilizer group $\mathbb{Z}_{n}$ for some $n\in\mathbb{N}$. There is a $G$-invariant tubular neighborhood of $U_{p}$ which is symplectomorphic to a trivial disc bundle $E\cong D\times U_{p}$ over $U_{p}$ on which $G$ acts nontrivially as rotations along fibers of $E$, fixing $U_{p}$ pointwise. Since $\Sigma$ is symplectic we may identify $E=D\times U_{p}$ with $D_{1}\times D_{2}\subset\mathbb{C}\times\mathbb{C}$ with $U_{p}=\{0\}\times D_{2}$ being the base disc centered at $p=(0,0)$, and $D_{1}\times\{pt\}$ as fibers, with $G$ acts $D_{1}\times D_{2}$ by $t\cdot(z_{1},z_{2})=(e^{\pm int}z_{1},z_{2})$, $t\in G\cong\mathbb{R}/2\pi\mathbb{Z}$.
Let $\gamma$ be a $G$-orbit in the disc fiber over $p$ and $D_{\gamma}\subset D_{1}\times\{0\}$ the fiber disc over $p$ bounding $\gamma$. Assume that $h(\Sigma)$ is a local minimum of $h$ then $h=\frac{n}{2}|z_{1}|^{2}+constant$ on $E$ and the tangent vector field of $\gamma$ is $X=n(-y_{1}\partial_{x_{1}}+x_{1}\partial_{y_{1}})$. Let $v:=\partial_{x_{2}}|_{\gamma}$. Then $L:=X\wedge v$ is a loop of Lagrangian planes along $\gamma$ with maslov index $\mu(L,D_{\gamma})=2n\neq 0$. On the other hand, if $h(\Sigma)$ is a local maximum of $h$ then $h=-\frac{n}{2}|z_{1}|^{2}+constant$ on $E$ and the tangent vector field of $\gamma$ is $X=-n(-y_{1}\partial_{x_{1}}+x_{1}\partial_{y_{1}})$. In this case $\mu(L,D_{\gamma})=-2n\neq 0$. Either way the Maslov condition is not met. So $Fix(G)$ does not contain any symplectic surface. We conclude that $Fix(G)$ is a discrete set of points. ∎
@@
Let $(M,\omega=d\lambda,G)$ be a connected, exact special Hamiltonian $S^{1}$-space and $\lambda$ a $G$-invariant primitive 1-form of $\omega$. Then

### v4

@@ 3 Some topological constraints imposed by Maslov condition
(ii) It follows from (i) that an equivariant version of Darboux Theorem [1] (see also [6] Theorem 22.2) applies to every isolated fixed point $p\in Fix(\Phi)$. Namely, for each isolated fixed point $p\in Fix(\Phi)$, there exists a $\Phi$-invariant chart $V$ centered at $p\in M$ and $\Phi$-equivariant with respect to a linear action of $\Phi$ on $\mathbb{R}^{4}$ so that $\omega|_{V}=\omega_{0}$ is the standard symplectic form $\omega_{0}=\sum_{k=1}^{2}dx_{k}\wedge dy_{k}$ on $\mathbb{R}^{4}$, and $f^{-1}\circ\Phi\circ f=\Theta\subset SU(2)$.
@@
For (iii), observe that $dh|_{p}=-\omega(X_{h}|_{p},\cdot)$ for $p\in M$. Since $\omega$ is nondegenerate, $\omega(X_{h}|_{p},\cdot)=0$ iff $X_{h}$ vanishes at $p$, i.e., iff $p\in Fix(\Phi)$, i.e., $Crit(h)=Fix(\Phi)$. By (ii) we may assume that $\Phi=\Theta$ and hence $h=\textsf{h}$ near each isolated critical point $p\in Fix(\Phi)$ of $h$, therefore all isolated critical points of $h$ are nondegenerate with Morse index $2$. ∎
@@
Forgetting the Maslov condition, 4-dimensional SHam1-spaces are by themselves (not necessarily compact) Hamiltonian $S^{1}$-spaces. Recall that 4-dimensional compact Hamiltonian $S^{1}$-spaces have been classified by Karshon [8]:
@@
    the complex projective plane $\mathbb{C}P^{2}$ or a Hirzebruch surface, with a symplectic form that comes from a Kähler form, if $M$ has at most one fixed surface (a symplectic sphere); or
@@
Recall that if $(M,\omega,\Phi)$ is a 4-dimensional compact Hamiltonian $S^{1}$-space then each connected component of $Fix(\Phi)$ is either a single point or a symplectic surface, and the maximum and minimum of the moment map $h$ is each attained on exactly one component of $Fix(\Phi)$ (see e.g. [8] Appendix A).
@@
A closed compact symplectic 4-manifold $(M,\omega)$ does not admit a SHam1-action.
@@
Assume in contrary that $(M,\omega)$ allows a special SHam1-action $\Phi:S^{1}\to Ham(M,\omega)$. We may assume that the action is effective. Let $h$ denote its moment map. Then $Fix(\Phi)=Crit(h)$. $M$ is compact, $h$ is bounded, and the sets $M_{-}:=\{p\in M\mid h(p)=\min h\}$ and $M_{+}:=\{p\in M\mid h(p)=\max h\}$ are both nonempty. Since every isolated fixed point of the $\Phi$-action has Morse index 2, the set $M_{-}$ contains no connected component with isolated fixed points, and neither does $M_{+}$. Then both $M_{-}$ and $M_{+}$ are connected compact closed symplectic surfaces and, following [8], $M$ can be obtained from a ruled manifold with two fixed surfaces and no interior fixed points by a sequence of equivalent blowups at fixed points that are not minima for the moment map of the ruled manifold.
Suppose that $M$ is a ruled manifold, then the orbit $\gamma_{q}:=Orb_{G}(q)$ of $q\not\in Fix(G)$ is contained some symplectic $S^{2}$-fiber $S$ of $M$. But $c_{1}(M)([S])=S\cdot S+e(S)=0+2\neq 0$, $c_{1}(M.\omega)\neq 0$. Indeed, let $F_{p}$ denote the $S^{2}$ fiber over $p\in M_{-}$. Fix any $q\in F_{p}\setminus(M_{+}\cup M_{-})$. $Orb_{\Phi}(q)$ divides $F_{p}\cong S^{2}$ into two discs $D_{-}\cup D_{+}$ with $D_{-}\cap M_{-}=\{p\}$. We may identity a $\Phi$-invariant symplectic Darboux neighborhood of $p\in M$ as $D^{\prime}\times D_{\epsilon}\subset\mathbb{C}_{z_{1}}\times\mathbb{C}_{z_{2}}$ so that $D_{-}\subset D^{\prime}\times\{0\}$, and $D_{\epsilon}$ is a coordinate chart of $F_{p}$ centered at $p$. Up to a choice of orientation of $\gamma_{q}$ we may assume that the tangent vector field along $\gamma_{q}$ is $X:=-y_{1}\partial_{x_{1}}+x_{1}\partial_{y_{1}}$. Let $v:=\partial_{x_{2}}|_{\gamma_{q}}$ be the vector field $\partial_{x_{2}}$ restricted to $\gamma_{q}$. Then $L:=X\wedge v$ is a Lagrangian plane field along $\gamma_{q}$ with Maslov index $\mu(L,D_{-})=2$. On the other hand, if we choose $D_{+}$ as the disc bounded by $\gamma_{q}$, then the orientation of $\gamma_{q}$ as the boundary of $D_{+}$ is given by X-X, and the corresponding Maslov index becomes $\mu(L=(-X)\wedge v,D_{+})=-2$. Either way the Maslov condition is not satisfied. So $M$ cannot be a ruled manifold, nor can it be obtained from any ruled manifold by a sequence of equivalent blowups at fixed points that are not minima for the moment map of the ruled manifold. We conclude that a 4-dimensional compact symplectic manifold $(M,\omega)$ does not admit an effective special Hamiltonian $S^{1}$-action. ∎
@@
Let $(M,\omega)$ be a connected noncompact symplectic 4-manifold. Suppose that there is a special Hamiltonian $S^{1}$-action $\Phi:S^{1}\to Ham(M,\omega)$ on $M$ with contractible orbits and nonempty fixed point set $Fix(\Phi)$, then $Fix(\Phi)$ consists of isolated points.
@@
Let $\Phi:S^{1}\to Ham(M,\omega)$ be a semi-free SHam1-action on $(M,\omega)$. Denote by $X$ the Hamiltonian vector field on $M$ which generates the $\Phi$-action, and $h:M\to\mathbb{R}$ the corresponding moment map defined by $\omega(X,\cdot)=-dh$. By assumption all $\Phi$-orbits are contractable and $Fix(\Phi)$ is not empty. It is known that every connected component of $Fix(\Phi)$ is either a single point or a symplectic surface ([8], Appendix A). If $Fix(\Phi)$ contains a connected symplectic surface $\Sigma$ then $h$ is constant on $\Sigma$ and $h(\Sigma)$ is a local extremum of $h$.
Fix a point $p\in\Sigma$ and an open neighborhood $U_{p}\subset\Sigma$ of $p$ so that $\Phi$ acts nontrivially near $U_{p}$ except on $U_{p}$. There is a $\Phi$-invariant tubular neighborhood of $U_{p}$ which is symplectomorphic to a trivial disc bundle $E\cong D\times U_{p}$ over $U_{p}$ on which $\Phi$ acts nontrivially as rotations along fibers of $E$, fixing $U_{p}$ pointwise. Since $\Sigma$ is symplectic we may identify $E=D\times U_{p}$ with $D_{1}\times D_{2}\subset\mathbb{C}\times\mathbb{C}$ with $U_{p}=\{0\}\times D_{2}$ being the base disc centered at $p=(0,0)$, and $D_{1}\times\{pt\}$ as fibers, with $\Phi$ acts $D_{1}\times D_{2}$ by $t\cdot(z_{1},z_{2})=(e^{\pm int}z_{1},z_{2})$, $t\in\Phi\cong\mathbb{R}/2\pi\mathbb{Z}$.
Let $\gamma$ be a $\Phi$-orbit in the disc fiber over $p$ and $D_{\gamma}\subset D_{1}\times\{0\}$ the fiber disc over $p$ bounding $\gamma$. Assume that $h(\Sigma)$ is a local minimum of $h$ then $h=\frac{1}{2}|z_{1}|^{2}+constant$ on $E$ and the tangent vector field of $\gamma$ is $X=-y_{1}\partial_{x_{1}}+x_{1}\partial_{y_{1}}$. Let $v:=\partial_{x_{2}}|_{\gamma}$. Then $L:=X\wedge v$ is a loop of Lagrangian planes along $\gamma$ with maslov index $\mu(L,D_{\gamma})=2\neq 0$. On the other hand, if $h(\Sigma)$ is a local maximum of $h$ then $h=-\frac{1}{2}|z_{1}|^{2}+constant$ on $E$ and the tangent vector field of $\gamma$ is $X=y_{1}\partial_{x_{1}}-x_{1}\partial_{y_{1}}$. In this case $\mu(L,D_{\gamma})=-2\neq 0$. Either way the Maslov condition is not met. So $Fix(\Phi)$ does not contain any symplectic surface. We conclude that $Fix(\Phi)$ is a discrete set of points. ∎
@@
Let $(M,\omega=d\lambda,\Phi)$ be a connected, exact, simple SHam1-space, and let $\lambda$ be a $\Phi$-invariant primitive 1-form of $\omega$. Then $Fix(\Phi)\subset h^{-1}(0)$ up to addition of a constant to $h$,
Proof.
Let $X$ be the Hamiltonian vector field of the moment map $h$ associated to the $\Phi$-action. Observe that
(2) $dh=d\lambda(X_{h},\cdot)=L_{X_{h}}\lambda-d(\lambda(X_{h}))=-d(\lambda(X_{h}))$
with $L_{X_{h}}\lambda=0$ since $\lambda$ is $\Phi$-invariant, so $\lambda(X_{h})-h$ is a constant. Up to an addition of a constant to $h$ we may assume that
(3) $h=\lambda(X_{h})\quad\text{and hence \quad$\lambda(X_{h})=0$ on $h^{-1}(0)$},$
then
$Fix(\Phi)\subset h^{-1}(0)$
since $\lambda(X_{h})$ is constant on level sets of $h$. ∎
4 Simply-connected Exact Simple SHam1-spaces
The rest of this paper will focus on the classification of connected exact simple SHam1-spaces. We assume that $Fix(\Phi)$ is a nonempty finite set as before. Let
$M_{c}:=\{h^{-1}(c)\},\quad\EuFrak{r}_{c}=\EuFrak{r}|_{h^{-1}(c)}:h^{-1}(c)/\Phi,\quad c\in\mathbb{R}.$
Proposition 4.1.
Let $(M,\omega=d\lambda,\Phi)$ be a connected, 1-connected exact simple SHam1-space with $c_{1}(M,\omega)=0$, $\lambda$ being $\Phi$-invariant, $\Phi$ acts semi-freely on $M$ with moment map $h:M\to\mathbb{R}$, and $Fix(\Phi)\subset h^{-1}(0)$ is a finite set. Then the followings are true.